OpenStudy (anonymous):

find vertex and roots for x^2+8x+12

OpenStudy (isaiah.feynman):

To find the vertex you can complete the square or use this formula..\[\left( -\frac{ b }{ 2a }, f \left( -\frac{ b }{ 2a } \right) \right).\]To find the roots, simply factor the quadratic as usual.

OpenStudy (anonymous):

ok